Simon Companys year-end balance sheets follow. Current Yr 1 Yr Ago 2 Yrs Ago $ At December 31 Assets Cash Accounts receivablComplete this question by entering your answers in the tabs below. Required 1A Required 1B Compute days sales uncollected. DCompute accounts receivable turnover. Accounts Receivable Turnover 1 Choose Denominator: Choose Numerator: Accounts ReceivablCompute inventory turnover. Inventory Turnover Choose Numerator: 1 Choose Denominator: = Inventory Turnover Inventory turnoveComplete this question by entering your answers in the tabs below. Required 4A Required 4B Compute days sales in inventory.
Answer #1
Day' s sales uncollected
Average Accounts Receivable / Net credit sales x Days = Day's sales uncollected
Current Year 75900 / 700837 x 365 = 40 Days
1 yr ago 56600 / 553048 x 365 = 37 Days
Accounts Receivable Turnover
Net sales / Average Accounts receivable = Accounts receivable Turnover
Current Year 700837 / 75900 = 9.23 times
1 yr ago 553048 / 56600 = 9.77 times
Inventory Turnover
Cost of goods sold / Average Inventory = Inventory Turnover
Current Year 427511 / 99250 = 4.31 times
1 yr ago 359481 / 69750 = 5.15 times
Day' s sales in Inventory
Average Inventory / Cost of goods sold x Days = Day's sales in Inventory
Current Year 99250 / 427511 x 365 = 85 Days
1 yr ago 69750 / 359481 x 365 = 71 Days